OpenStudy (gokuporter):

Sixty is what percent of 50? A. 30% B. 90% C. 110% D. 120%

OpenStudy (anonymous):

you do .60* 50 which is 30

OpenStudy (holly00d1248):

I agree 30%

OpenStudy (anonymous):

D. 120%

OpenStudy (gokuporter):

you took the test @collins-fortuchang

OpenStudy (anonymous):

yes

OpenStudy (gokuporter):

@collins-fortuchang can you help me

OpenStudy (anonymous):

sure

OpenStudy (gokuporter):

Kiara estimated that 45 customers bought fruit at her farm stand one day. The actual number of customers who bought fruit was 41. What was the percent error in Kiara’s estimate? Round to the nearest tenth of a percent, if needed. A. 2% B. 4% C. 8.9%. D. 9.8%

OpenStudy (gokuporter):

do you know this? @collins-fortuchang

OpenStudy (anonymous):

i have to got get my journal just post all the questions

OpenStudy (gokuporter):

A $5000 investment earns $1575 simple interest in three years. What is the annual interest rate? A. 1.05% B. 3.15% C. 10.5% D. 31.5%

OpenStudy (gokuporter):

Lucas bought some plumbing equipment for $37.38. The price he paid included a sales tax of 5%. What was the price of the plumbing equipment before the tax was added? A. $35.60 B. $36.33 C. $37.33 D. $39.25

OpenStudy (gokuporter):

The price of a bag of oranges is directly proportional to the number of oranges packaged in the bag. The table shows this relationship. What is the constant of proportionality?

OpenStudy (gokuporter):

The value of y varies directly with the value of x, and y = –21 when x = 3. Which equations represent the relationship between y and x? Choose exactly two answers that are correct.

OpenStudy (gokuporter):

@collins-fortuchang you done yet?

OpenStudy (anonymous):

for the question with kiara if u round it u get 9.8 for number 2 in 1.05 number 39.25 for the next one use k=y/x fro the final one y=7

OpenStudy (gokuporter):

k=y/x?

OpenStudy (gokuporter):

@collins-fortuchang I don't get that one

OpenStudy (gokuporter):

@collins-fortuchang you there?

OpenStudy (gokuporter):

http://prntscr.com/66w1gf is this correct?

OpenStudy (anonymous):

k is for constant of proportionality so u divide the y by the x

OpenStudy (anonymous):

yes

OpenStudy (gokuporter):

@collins-fortuchang http://prntscr.com/66w3oq is this correct

OpenStudy (anonymous):

yes

OpenStudy (gokuporter):

@collins-fortuchang http://prntscr.com/66w6bh is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

divide 7.20 by 4 then multiply by 9